coach reed is ordering for his 18 member soccer team. He will be ordering a $1.35 drink and a hamburger for each member of his team. The bill without tax comes to $54.72. Write and solve an equation to show the cost of a hamburger
Answer by stanbon(75887)

coach reed is ordering for his 18 member soccer team. He will be ordering a $1.35 drink and a hamburger for each member of his team. The bill without tax comes to $54.72. Write and solve an equation to show the cost of a hamburger
----
d + h = 18
Value Eq:: 18[1.35 + h] = 54.72
1.35+h = 3.04
h = $1.69 (cost of one hamburger)
----
Cheers,
Stan H.
---------